# On the Approximation Power of SiLU Networks: Exponential Rates and Depth Efficiency

### v1

@@ A.6 Proof of Theorem 8
$\left|f(\mathbf{x})-f_{\varepsilon}(\mathbf{x})\right|$ $=\left|\sum_{\mathbf{j}}\psi_{\mathbf{j}}(\mathbf{x})\left(f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\right)\right|$
$\leq\sum_{\mathbf{j}\in C_{\mathbf{j}}}\left|\left(f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\right)\right|,\qquad\text{since}\ \psi_{\mathbf{j}}(\mathbf{x})\leq 1$
$\leq 2^{d}\max_{\mathbf{j}\in C_{\mathbf{j}}}\left|\left(f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\right)\right|$
$\leq 2^{d}\frac{h^{n}}{n!2^{n}}\max_{|\bm{\alpha}|=n}\text{ess sup}_{\mathbf{x}\in[-B,B]^{d}}\left|D^{\bm{\alpha}}f(\mathbf{x})\right|$
$\leq 2^{d}\frac{h^{n}}{n!2^{n}}$
The last three inequalities hold because respectively: in each direction $\mathbf{j}$, there are at most $2$ cubes bounding $C_{\mathbf{j}}$; for $\mathbf{x}\in C_{\mathbf{j}}$, we have $|x_{i}-c_{j_{i}}|\leq\frac{h}{2}$ for all $i$, hence
$|\mathbf{x}-\mathbf{c}_{\mathbf{j}}|^{\bm{\alpha}}=\prod_{i=1}^{d}|x_{i}-c_{j_{i}}|^{\alpha_{i}}\leq\left(\frac{h}{2}\right)^{|\bm{\alpha}|}=\left(\frac{h}{2}\right)^{n},$
and $\text{ess sup}_{\mathbf{x}\in[-B,B]^{d}}\left|D^{\bm{\alpha}}f(\mathbf{x})\right|\leq 1.$
We want $\|f-f_{\varepsilon}\|_{\infty}\leq\frac{\varepsilon}{2}$. It suffices that
$2^{d}\frac{h^{n}}{n!2^{n}}\leq\frac{\varepsilon}{2}\quad\Rightarrow\quad h\leq 2\left(\frac{(n!)\varepsilon}{2^{d+1}}\right)^{1/n}.$
Since $h=\frac{2B}{M}$, we choose:
@@
$f_{\varepsilon}(\mathbf{x})=\sum_{\mathbf{j}\in M_{d}}\sum_{|\bm{\alpha}|\leq n-1}\underbrace{\frac{D^{\bm{\alpha}}f(\mathbf{c}_{\mathbf{j}})}{\bm{\alpha}!}}_{C_{\mathbf{j},\bm{\alpha}}}\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}},\qquad\text{with}\ C_{\mathbf{j},\bm{\alpha}}\leq 1$
  • According to Theorem 3, we approximate each term $\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}=\prod_{i=1}^{d}(x_{i}-c_{j_{i}})^{\alpha_{i}}$, a product of $d+n$ factors
    $\mathbf{1}_{[\alpha_{j_{1}},\beta_{j_{1}}]}(x_{1})\times\mathbf{1}_{[\alpha_{j_{2}},\beta_{j_{2}}]}(x_{2})\times\cdots\times\mathbf{1}_{[\alpha_{j_{d}},\beta_{j_{d}}]}(x_{d})\times(x_{1}-c_{j_{1}})\times\cdots\times(x_{2}-c_{j_{2}})\times\cdots\times(x_{d}-c_{j_{d}})\times\cdots$
@@
    $\widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})=\widetilde{M}_{k}\left(\phi_{\alpha_{j_{1}},\beta_{j_{1}},\tau,\kappa}(x_{1}),\widetilde{M}_{k}\left(\phi_{\alpha_{j_{2}},\beta_{j_{2}},\tau,\kappa}(x_{2}),\ldots,\widetilde{M}_{k}\left((x_{1}-c_{j_{1}}),(x_{1}-c_{j_{1}})\right)\right),\ldots\right)$
@@
    $\leq\begin{cases}\eta+\gamma_{i},\ \text{If}\ \ |u_{i+1}|=|\phi_{\alpha_{j_{i+1}},\beta_{j_{i+1}},\tau,\kappa}(x_{i+1})|\leq 1\\ \eta+2\gamma_{i},\ \text{If}\ \ |u_{i+1}|=|x_{i+1}-c_{j_{i}}|\leq B\leq 2,\end{cases}$
    During successive approximation, the corresponding error evolves linearly:
    $\varepsilon_{\text{m}}=\eta\sum_{j=n}^{d+n-2}1^{j}+\eta\sum_{j=0}^{n-2}2^{j}\leq d\eta+\eta(2^{n-2}-1).$
@@
$k\geq\max\left\{\left\lceil\frac{\ln C_{2}-\ln\eta}{2\ln\omega_{2}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{2}}\right\lceil\right\}=t_{0}(n,d)(\ln 1/\varepsilon+t_{1}(n,d))$
@@
We construct the network by partitioning $B,B]^{d}$ into $M^{d}=O(\varepsilon^{-d/n})$ cubes through the choice $M=\lceil B(2^{d+1}/(n!\varepsilon))^{1/n}\rceil$, and locally approximate $f$ by its Taylor polynomial $T_{\mathbf{j}}$ of order n1n-1 at each center $\mathbf{c}_{\mathbf{j}}$, then define the global approximant $f_{\varepsilon}(\mathbf{x})=\sum_{\mathbf{j}}\psi_{\mathbf{j}}(\mathbf{x})T_{\mathbf{j}}(\mathbf{x})$ where $\psi_{\mathbf{j}}$ is the indicator function of cube $C_{\mathbf{j}}$, ensuring $\|f-f_{\varepsilon}\|_{\infty}\leq\varepsilon/2$. Each term $\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\alpha}$ is a product of at most d+n1d+n-1 factors, approximated by a network $\widetilde{G}_{\mathbf{j},\alpha}$ using an approximate multiplication $\widetilde{M}_{k}$ of depth 1 and width 2, with per-multiplication error $\eta$ chosen such that $\eta(2^{d+n-1}-1)\leq\varepsilon/(2^{d+1}d^{n})$ by taking $k\sim\ln(1/\varepsilon)$. The $\mathcal{O}(\varepsilon^{-d/n})$ terms $\widetilde{G}_{\mathbf{j},\alpha}$ are computed in parallel over $\mathcal{O}(1)$ layers with width $\mathcal{O}(\varepsilon^{-d/n})$, then summed linearly via the addition of one linear layer, yielding a network $\widetilde{f}_{\varepsilon}$ with total depth $\mathcal{O}(1)+1=c(d,n,B)$ and size $\mathcal{O}(\varepsilon^{-d/n})$ satisfying $\|f-\widetilde{f}_{\varepsilon}\|_{\infty}\leq\varepsilon$. ∎
@@ Appendix B Finite Difference Operator $\Delta$
The finite difference operator $\Delta$ is define in [jordan1965calculus].
@@
Lemme 2.
$\Delta x^{\underline{k}}=kx^{\underline{k-1}}$ and $\Delta^{m}x^{\underline{k}}=k(k-1)\dots(k-m+1)x^{\underline{k-m}}\quad\text{if }k\geq m$
@@
$\Delta x^{\underline{k}}=(x+1)^{\underline{k}}-x^{\underline{k}}$
Now $(x+1)^{\underline{k}}=(x+1)x(x-1)\dots(x-k+2)$ and $x^{\underline{k}}=x(x-1)\dots(x-k+1)$.
Factoring by $x(x-1)\dots(x-k+2)$, the difference is written
$\Delta x^{\underline{k}}=[(x+1)-(x-k+1)]\cdot x(x-1)\dots(x-k+2)=k\cdot x^{\underline{k-1}}$
We prove the second equality by induction. for $m=1,$ the second equality is true because it is equivalent to the first. Assume that
$\Delta^{m}x^{\underline{k}}=k(k-1)\dots(k-m+1)\,x^{\underline{k-m}},\quad k\geq m.$
is true. Then
$\Delta^{m+1}x^{\underline{k}}=\Delta\left[k(k-1)\dots(k-m+1)\,x^{\underline{k-m}}\right].$
The constant factor $k^{\underline{m}}:=k(k-1)\dots(k-m+1)$ is factored out of the operator $\Delta$:
$\Delta^{m+1}x^{\underline{k}}=k^{\underline{m}}\cdot\Delta x^{\underline{k-m}}.$
But
$\Delta x^{\underline{k-m}}=(k-m)\,x^{\underline{k-m-1}}.$
Therefore
$\Delta^{m+1}x^{\underline{k}}=k^{\underline{m}}\cdot(k-m)\,x^{\underline{k-m-1}}=k^{\underline{m+1}}\,x^{\underline{k-(m+1)}}.$
Appendix C An upper bound of $|\sigma^{(n)}(a)|$
The function $\sigma$ is analytic in $\mathbb{C}$, in particular in a disk $D(a,r)$,
$\sigma^{(n)}(a)=\frac{n!}{2\pi i}\int_{|x-a|=r}\frac{\sigma(x)}{(x-a)^{n+1}}dx.$
An upper bound gives
$|\sigma^{(n)}(a)|=\left|\frac{n!}{2\pi i}\int_{|x-a|=r}\frac{\sigma(x)}{(x-a)^{n+1}},dx\right|$ $\leq\frac{n!}{2\pi}\int_{|x-a|=r}\left|\frac{\sigma(x)}{(x-a)^{n+1}}\right||dx|$
$\leq\frac{n!}{2\pi r^{n+1}}\left(\max_{|x-a|=r}|\sigma(x)|\right)\int_{|x-a|=r}|dx|$
$\leq\frac{n!}{r^{n}}\left(\max_{|x-a|=r}|\sigma(x)|\right)$
The poles of $\sigma(x)$ are solutions of $1+e^{-x}=0$, i.e., $e^{-x}=-1$, so $x_{k}=i\pi+2ik\pi$, $k\in\mathbb{Z}$. The distance from $a\in\mathbb{R}$ to the nearest poles is $d_{a,\pi}=\sqrt{a^{2}+\pi^{2}}\geq\pi$. If $r<d_{a,\pi}$, then in the circle $\mathcal{C}_{r}=\{x\in\mathbb{C},|x-a|=r\}$, the function $x\mapsto|1+e^{-x}|$ is a continuous function (because there are no poles on the circle if $r<d_{a,\pi}$), so it attains a minimum $\min_{|x-a|=r}|1+e^{-x}|=\delta(a,r)>0$ on this compact set. Then for all $x$ on this circle,
$|1+e^{-x}|\geq\delta(a,r)\quad\Rightarrow\quad\frac{1}{|1+e^{-x}|}\leq\frac{1}{\delta(a,r)}.$
Consequently,
$|\sigma(x)|\leq\frac{1}{\delta(a,r)}\quad\text{for }x\ \text{such that}|x-a|=r,$
with $r<d_{a,\pi}.$ and it follows the upper bound
$|\sigma^{(n)}(a)|\leq C_{a}n!,$ (13)
where $C_{a}=\frac{1}{\delta(a,r)r^{n}}.$

### v2

@@ 3.5 Approximation of Continuous Functions
$\begin{split}|f(\mathbf{x})-f_{\varepsilon}(\mathbf{x})|&=\Bigl|\sum_{\mathbf{j}}\psi_{\mathbf{j}}(\mathbf{x})\bigl(f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\bigr)\Bigr|\\ &\leq\sum_{\mathbf{j}\in C_{\mathbf{j}}}\bigl|f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\bigr|,\quad(\psi_{\mathbf{j}}(\mathbf{x})\leq 1)\\ &\leq 2^{d}\max_{\mathbf{j}\in C_{\mathbf{j}}}\bigl|f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\bigr|\\ &\leq 2^{d}\frac{h^{n}}{n!2^{n}}\max_{|\bm{\alpha}|=n}\operatorname*{ess\,sup}_{\mathbf{x}\in[-B,B]^{d}}|D^{\bm{\alpha}}f(\mathbf{x})|\\ &\leq 2^{d}\frac{h^{n}}{n!2^{n}}\end{split}$
The last three inequalities hold because respectively: At each point, at most $2d$ functions $\psi_{\mathbf{j}}$ are nonzero; for $\mathbf{x}\in C_{\mathbf{j}}$, we have $|x_{i}-c_{j_{i}}|\leq\frac{h}{2}$ for all $i$, and $\text{ess sup}_{\mathbf{x}\in[-B,B]^{d}}\left|D^{\bm{\alpha}}f(\mathbf{x})\right|\leq 1.$
We want $\|f-f_{\varepsilon}\|_{\infty}\leq\frac{\varepsilon}{2}$. It suffices to choose:
@@
$f_{\varepsilon}(\mathbf{x})=\sum_{\mathbf{j}\in M_{d}}\sum_{|\bm{\alpha}|\leq n-1}\underbrace{\frac{D^{\bm{\alpha}}f(\mathbf{c}_{\mathbf{j}})}{\bm{\alpha}!}}_{C_{\mathbf{j},\bm{\alpha}}}\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}},$
with $C_{\mathbf{j},\bm{\alpha}}\leq 1$. According to Theorem 3, we approximate each term $\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}=\prod_{i=1}^{d}\mathbf{1}_{\mathbf{C}_{j_{i}}}(x_{i})\prod_{i=1}^{d}(x_{i}-c_{j_{i}})^{\alpha_{i}}$, a product of $d+n$ factors
$\mathbf{1}_{\mathbf{C}_{j_{1}}}(x_{1})\times\mathbf{1}_{\mathbf{C}_{j_{2}}}(x_{2})\times\cdots\times\mathbf{1}_{\mathbf{C}_{j_{d}}}(x_{d})\times(x_{1}-c_{j_{1}})\times\cdots$
@@
$\widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})=\widetilde{M}_{k}\Bigl(\phi_{a_{j_{1}},b_{j_{1}},\tau,\kappa}(x_{1}),\widetilde{M}_{k}\bigl(\phi_{a_{j_{2}},b_{j_{2}},\tau,\kappa}(x_{2}),\ldots,\\ \widetilde{M}_{k}\bigl((x_{1}-c_{j_{1}}),\ldots\bigr)\bigr)\Bigr)$
@@
$\leq\begin{cases}\eta+\gamma_{i},\ \text{If}\ \ |u_{i+1}|=|\phi_{\alpha_{j_{i+1}},\beta_{j_{i+1}},\tau,\kappa}(x_{i+1})|\leq 1\\ \eta+2\gamma_{i},\ \text{If}\ \ |u_{i+1}|=|x_{i+1}-c_{j_{i}}|\leq 2B\leq 2,\end{cases}$
@@
$k\geq\left\lceil\frac{\ln C_{2}-\ln\eta}{2\ln\omega_{2}}\right\rceil=\mathcal{O}(\ln 1/\varepsilon)$
@@
The $\mathcal{O}(\varepsilon^{-d/n})$ terms $\widetilde{G}_{\mathbf{j},\alpha}$ are computed in parallel over $\mathcal{O}(1)$ layers with width $\mathcal{O}(\varepsilon^{-d/n})$, then summed linearly via the addition of one linear layer, yielding a network $\widetilde{f}_{\varepsilon}$ with total depth $\mathcal{O}(1)=c(d,n,B)$ and size $\mathcal{O}(\varepsilon^{-d/n})$ satisfying $\|f-\widetilde{f}_{\varepsilon}\|_{\infty}\leq\varepsilon$.
4 Discussions and experimental results
4.1 Parameterization of the Construction
Figure 10: Error evolution as a function of $\beta$ and $k$.
Figure 11: Error evolution as a function of $\beta$ and $a$.
Scale Parameter $\beta$
The parameter $\beta$ determines the compression rate in the recursive construction $Q_{k}(x)=\beta^{2k}f_{a}(\beta^{-k}x)$. Its optimal value results from a trade-off between convergence speed and numerical stability. A systematic analysis reveals that values that are too small (e.g., $0.01\leq\beta\leq 0.5$, Fig. 10) theoretically accelerate convergence. These values $\beta of$fer a good compromise, allowing for low values of $k$ to achieve maximum reconstruction errors on the order of $10^{-7}$.
Recursion Parameter $k$
The parameter $k$, less influenced by $a$, controls the accuracy of the approximation according to the relation $Q_{k}(x)=\beta^{2k}f_{a}(\beta^{-k}x)$. Experimentally, we observe that $k=3$ is sufficient to achieve an accuracy of $10^{-6}$ in the approximation of the function $x^{2}$ (Fig. 12), while $k=5$ allows reaching $10^{-15}$ (Fig. 10).
Symmetrization Shift $a$
The parameter $a$ in the function $g_{a}(x)=(x+a)\sigma(x+a)+(-x+a)\sigma(-x+a)-2a\sigma(a)$ influences the dominant coefficient $K(a)=2\sigma^{\prime}(a)+a\sigma^{\prime\prime}(a)$. Numerical analysis shows that the Taylor expansion remains well‑conditioned when centered at $a=0$ or $a=1$, provided the function and the evaluation points are appropriately scaled (Fig. 11).
4.2 Approximation of Continuous Functions
Our experiments (Figs. 13 - 16) implemented networks of by theorems 6, 7 and 8 to approximate the following functions:
$T_{1}(x)=\log(7+x)\cos(x^{3}),\ T_{2}(x)=\frac{1}{1+e^{-x}},\ T_{3}(x)=(x,\sin(x),\cos(x)),\ T_{4}(x,y)=\frac{\sin(\pi x)\cos(\pi x)}{2\pi}.$
Figure 13: Approximation of the function $\log(7+x)\cos(x^{3})$ by the NN of Theorem 6 (in left) and Theorem 7 (in right).
Figure 14: Approximation of the sigmoid function $\frac{1}{1+e^{-x}}$ by the NN of Theorem 6 (in left) and Theorem 7 (in right).
Figure 16: Approximation of the function $(x,y)\mapsto\frac{\sin(\pi x)\cos(\pi x)}{2\pi}$ (in left) by theorem 8 (in right).
Based on some experimental results (Figs. 13 and 14), we observe good accuracy in the approximation by neural networks with SiLU activation functions via the step functions. The theoretical construction in proof of theorem 8 is experimentally validated with $B=1$, $n=5$, $M=3$ and $2000$ training points, confirming the practical efficiency of SiLU networks for Sobolev function approximation (Fig. 16).
Performance Analysis
Experimental implementation (Figs. 17 and 18) of the architectures in Theorems 1 and 3 for approximating $x^{2}$ and $xy$ highlights the crucial role of parameter initialization. SiLU networks initialized near the analytical solution converge via backpropagation to values close to the theoretical optimum, whereas ReLU networks, whith same architecture, often become trapped in suboptimal minima, yielding less accurate approximations. Our work shows that by allowing weights to scale with precision requirements, one can achieve constant-depth networks for Sobolev approximation. While the total bit complexity remains $\mathcal{O}(\varepsilon^{-d/n}\ln(1/\varepsilon))$ as in Yarotsky’s work, the architectural depth is reduced from logarithmic to constant.
Figure 18: The same observations hold for the network approximating the product $\cos x\cdot\sin x\log(x^{2})$ using the architecture from Theorem 2.
5 Conclusion
This work develops constructive approximation methods for SiLU networks, showing that with precision-tunable activation parameters, one can build constant-depth networks for Sobolev functions using $\mathcal{O}(\varepsilon^{-d/n})$ parameters. Our explicit constructions exhibit exponential convergence for basic operations and reveal parameter-efficient structures for polynomial approximation. Compared to logarithmic-depth ReLU networks, our approach trades architectural depth for increased weight precision, offering an alternative efficiency trade-off. These results advance the understanding of parametric activation networks and open up applicative perspectives in scientific computing where constant-depth architectures may prove advantageous.
Acknowledgments
This work was supported by the Agence Française de Développement (AFD) through the Agence Nationale de la Recherche (ANR-21-PEA2-0007), under the Partenariats with "l’Enseignement supérieur Africain" (PEA) program.
Appendix A An upper bound of $|\sigma^{(n)}(a)|$
The function $\sigma(x)=(1+e^{-x})^{-1}$ is analytic in $\mathbb{C}$. For any $a\in\mathbb{R}$ and $0<r<\sqrt{a^{2}+\pi^{2}}$, Cauchy’s formula gives:
$|\sigma^{(n)}(a)|\leq\frac{n!}{r^{n}}\max_{|x-a|=r}|\sigma(x)|.$
The poles of $\sigma$ are at $x_{k}=i\pi(1+2k)$, $k\in\mathbb{Z}$, hence the distance from $a$ to the nearest pole is $d(a)=\sqrt{a^{2}+\pi^{2}}\geq\pi$. Choosing $r=\pi/2$ (which satisfies $r<d(a)$ for all $a\in\mathbb{R}$), we have:
$\delta(a):=\min_{|x-a|=\pi/2}|1+e^{-x}|>0,$
and consequently
$\max_{|x-a|=\pi/2}|\sigma(x)|\leq\frac{1}{\delta(a)}.$
Therefore,
$|\sigma^{(n)}(a)|\leq\frac{n!}{(\pi/2)^{n}}\cdot\frac{1}{\delta(a)}=n!\cdot\left(\frac{2}{\pi}\right)^{n}\cdot C(a),$
where $C(a)=1/\delta(a)$ is independent of $n$. In particular, there exists a constant $C^{\prime}_{a}$ such that
$\left|\frac{\sigma^{(n)}(a)}{n!}\right|\leq C^{\prime}_{a}\left(\frac{2}{\pi}\right)^{n}.$
@@ Appendix B Finite Difference Operator $\Delta$
The finite difference operator $\Delta$ is define in [10].
@@
Lemme 2 (Differences of falling factorials).
For $k\geq m$,
$\Delta x^{\underline{k}}=k\,x^{\underline{k-1}},\qquad\Delta^{\,m}x^{\underline{k}}=k^{\underline{m}}\,x^{\underline{k-m}},$
where $k^{\underline{m}}=k(k-1)\dots(k-m+1)$ is the falling factorial of $k$.
@@
Using $\Delta x^{\underline{k}}=(x+1)^{\underline{k}}-x^{\underline{k}}$ and factoring $x(x-1)\dots(x-k+2)$ gives
$\Delta x^{\underline{k}}=\big[(x+1)-(x-k+1)\big]\,x(x-1)\dots(x-k+2)=k\,x^{\underline{k-1}}.$
The formula for $\Delta^{\,m}$ follows by induction on $m$: assuming it holds for $m$,
$\Delta^{\,m+1}x^{\underline{k}}=\Delta\bigl[k^{\underline{m}}\,x^{\underline{k-m}}\bigr]=k^{\underline{m}}\,(k-m)\,x^{\underline{k-m-1}}=k^{\underline{m+1}}\,x^{\underline{k-(m+1)}}.\qed$